OpenStudy (anonymous):

Find the equation in standard form which passes through (3, –6) and (–2, –1). You must show all work.

OpenStudy (anonymous):

(3,-6)-(-2,-1) To calculate the equation of the line, use the y=mx+b format. m=(-1-(-6))/(-2-(3)) y=mx+b (-6)=(-1)*(3)+b b=-3 y=-x-3

OpenStudy (anonymous):

wait is that the answer or do i need to solve it @Luis_Rivera

OpenStudy (anonymous):

what else do you need. its solved and answrd !

OpenStudy (anonymous):

no im sorry i thought their was more thank you @Luis_Rivera

OpenStudy (anonymous):

kk :)

OpenStudy (anonymous):

@MilaCoolNess plz close the question :)

OpenStudy (anonymous):

@Aperogalics im reviewing it ill close it when im done

OpenStudy (anonymous):

k :)